if (a.a.a) + 4(a) - 8 = 0 . Then what is value of (a.a.a.a.a.a.a) + 64(a.a)

Grade:8

2 Answers

jithin p nair
12 years ago

(a.a.a.a.a.a.a)+64(a.a)   =    a(a^6+64a)

but from the given condition we get:

a^3= (8-4a)

thus a(a^6+64a)=a((8-4a)^2+64a)

 

                        =a(64+16a^2-64a+64a)

 

                         =a(64+16a^2)

 

                        = 64a+16a^3

 

                        = 64a +16(8-4a)

 

                       =64a+128-64a

 

                       =128

Gyy
one year ago

(a.a.a.a.a.a.a)+64(a.a)   =    a(a^6+64a)

but from the given condition we get:

a^3= (8-4a)

thus a(a^6+64a)=a((8-4a)^2+64a)

 

                        =a(64+16a^2-64a+64a)

 

                         =a(64+16a^2)

 

                        = 64a+16a^3

 

                        = 64a +16(8-4a)

 

                       =64a+128-64a

 

                       =128